·        A vertical spring has a force constant of 125 N/m. by how much will the spring be compressed if a mass of 0.525kg placed on top of it?


1
Expert's answer
2021-02-22T10:21:12-0500

We can find the compression of the spring from the Hooke's law:


"F=kx,""mg=kx,""x=\dfrac{mg}{k}=\dfrac{0.525\ kg\cdot 9.8\ \dfrac{m}{s^2}}{125\ \dfrac{N}{m}}=0.041\ m."

Answer:

"x=0.041\ m."